# Probing quantum complexity via universal saturation of stabilizer entropies

### v1

@@ V Random Hamiltonian evolution
Finally, as we show in Appendix A, the Pauli spectrum and SRE of the GUE evolution matches closely the dynamics of the random basis evolution of Sec. IV. We also observe that the SREs for both models match. Thus, we argue that the scale-invariant behavior that we shown analytically in Sec. IV for random bases evolution also emerges for the evolution with random Hamiltonians.
VI Discussion
@@
We studied $\alpha$-SREs for random Clifford circuits doped with T-gates and random time evolution where we demonstrated the connection of Rényi index $\alpha$ to different aspects of complexity of quantum states.
@@
We find that the SRE converges to the maximum at a critical T-gate density $q_{\text{c},\alpha}$ and time $t_{\text{c},\alpha}$ in the thermodynamic limit. We determine the transition exactly for $\alpha=2$, while for general $\alpha$ we determine the convergence using heuristic models of the Pauli spectrum. For $\alpha=2$, we observe universal behavior around the critical point where the derivative of the SRE can be rescaled onto a single curve for all $N$. This hints that the saturation transition is connected to phase transitions, where universal behavior is commonly found, for example for the transition between different phases of quantum many-body system [35] or at complexity transitions in classical and quantum algorithms [39].
@@
We observe that for $\alpha>1$, the SRE is similar to $D_{\text{min}}$, which is the distance to the closest stabilizer state [16]. In fact, we show that the previously proven lower bound [19] is indeed tight for random evolutions Eq. (24) which we numerically confirm in Appendix C. Note that numeric evidence shows that SREs for $\alpha>1$ also provide an $N$-independent upper bound to $D_{\text{min}}$ [16]. As such, we argue that SREs with $\alpha>1$ probe the closeness to the nearest stabilizer state. SREs with $\alpha=2$ also have an operational meaning: They give a lower bound on fidelity estimation [40, 23]: Given a target state |$\ket{\psi}$, one can estimate the fidelity with actual state $\rho$ by just measuring Pauli expectation values [40]. The number of samples $m$ to estimate the fidelity is lower bounded as $m\gtrsim\exp(M_{2}(\ket{\psi}))$ (see also Appendix D).
In contrast, SREs with $\alpha<1$ (especially $\alpha=1/2$) show behavior similar to the log-free robustness of magic LR [2, 41, 4] or max-relative entropy of magic [9, 4]. They respectively relate to the negativity of the mixture of stabilizer states, or the number of superpositions of stabilizer states needed to represent a given state. LR has been used to estimate fault-tolerant state preparation complexity and relates to the complexity of Clifford based simulation algorithms. These algorithms simulate quantum circuits as Clifford circuits injected with nonstabilizer gates, where the simulation complexity commonly increase exponentially with the number of nonstabilizer gates [3]. In fact, $M_{1/2}$ has been used as a proxy for LR to evaluate simulation complexity of Clifford-based simulation algorithms [41]. Further, we find that the lower bound $\text{LR}\geq\frac{1}{2}M_{1/2}$ Eq. (8) is tight for random evolution (see Appendix C). Additionally, $M_{0}$ is a lower bound to the stabilizer nullity $\nu\geq M_{0}$, which characterizes the complexity of Clifford-based learning algorithms [42, 43, 44].
We study the dynamics of two models: First, for Clifford gates injected with T-gates, SREs converge for all $\alpha$ to their maximum $M_{\alpha}\sim N$ at a linear number of T-gates. This is because each T-gate affects only a discrete subset of the Pauli spectrum. We numerically find that $D_{\text{min}}$ and LR appear to show this behavior as well. This implies that fidelity with the closest stabilizer state and classical cost of simulation with Clifford+T correlate. In particular, for $N_{\text{T}}=\text{const}$ and thus $M_{\alpha}=\text{const}$, one can efficiently simulate and learn the state [42, 43, 44], as well as estimate the fidelity [23]. In contrast, for $N_{\text{T}}\sim N$ and thus $M_{\alpha}\sim N$ simulation, learning and fidelity estimation is unlikely to be efficient [45, 3, 5].
In contrast, SREs for random evolution shows widely different behavior depending on $\alpha$. This is because random evolution affects all Pauli strings even at short evolution times. For $\alpha>1$, $M_{\alpha>1}$ grows as $\propto t^{2}$ and converges to its maximum at $t_{\text{c},\alpha>1}\sim\sqrt{N}$. For $t=\text{const}$, we have $M_{\alpha>1}=\text{const}$ and $D_{\text{min}}=\text{const}$, implying that random evolution is close to a stabilizer state in terms of fidelity. This also implies that a polynomial number of Pauli measurements can certify the fidelity of quantum states (see Appendix D).
$\alpha<1$ shows completely different behavior: For $\alpha=0$, $M_{0}$ is maximal for any $t>0$, which implies that stabilizer nullity $\nu$ is maximal, rendering known near-Clifford learning algorithms inefficient [42, 43, 44]. For $10<\alpha<$ 1, SREs saturate rapidly at constant evolution time $t_{\text{c},0<\alpha<1}\approx\frac{1}{2}$. Further, extensive $M_{\alpha<1}\sim N$ is reached already for $t\gtrsim 1/\text{poly}(N)$. This hints that simulating random dynamics with Clifford simulation algorithms becomes classically intractable already at $t\gtrsim 1/\text{poly}(N)$.
@@
Thus, we find that random evolution has a fundamental separation in nonstabilizerness complexity: Simulation with Clifford-based algorithms is hard for $t\gtrsim 1/\text{poly}(N)$, while certifying the fidelity by measuring its Pauli strings is sample efficient. In contrast, for Cliffords doped with T-gates, simulation and certification complexity of aforementioned algorithms correlates and becomes intractable for the same T-gate density $q\gtrsim\log(N)$.
@@
Finally, we observe that for both random evolution and Clifford+T model that the critical time and T-gate density is maximal for $\alpha=2$. This indicates that the $2$-SRE holds a special status. Coincidentally, for $\alpha<2$, the SRE is known not to be a monotone [19], while for $\alpha\geq 2$ it is a pure state monotone [18].
@@
Finally, we want to highlight the technical contributions of our work: We show heuristically that the Pauli spectrum of random Hamiltonian evolution can be approximated by two distinct peaks. With increasing time, the two peaks shift towards each other and eventually merge. This is exactly when the SRE becomes maximal. At last, we introduce a class of random evolution in Eq. (19), which can be seen as evolution in random Clifford bases. This evolution behaves very similar to random Hamiltonian evolution, where we observe numerically the same Pauli spectrum. It can be expressed as random Clifford circuits combined with small-angle single-qubit rotations. This allows us to compute its $2$-SRE analytically for all times $t$. The random Clifford bases evolution could serve as a model of random evolution with an exact circuit representation. We believe these results may be of independent interest.
Acknowledgements.
@@ Appendix A GUE evolution and random basis evolution
While we believe that both evolutions show similar behavior for polynomial times, we note that for very long times (on the scale of $t\sim 2^{N/2}$ ) both models likely show different behavior in terms of deep thermalization [37], as the GUE Hamiltonian evolution conserves energy while the other model does not. It has been noted that the ensemble of GUE evolutions forms an exact $k$-design at polynomial times, however stops being a $k$-design at exponential times. This behavior at long times is attributed to energy conservation of the evolution, which at long times leads to a dephasing due to the energy eigenvalues. For non-energy conserved dynamics this behavior at long times is not expected. However, this difference at exponential times is evident in the $k$-design properties, however it may not be evident in the Pauli spectrum and SRE [16].
@@ Appendix D State certification via Pauli measurements and SREs
A common task is state certification to check whether the prepared state $\rho$ is close to the ideal state |$\ket{\psi}$ that one actually wanted to prepare. For this task, Ref. [40] proposed a simple algorithm that only requires to measure Pauli strings of the actual state. First, note that one can decompose any state in terms of its Pauli strings, i.e. $\rho=2^{-N}\sum_{\sigma\in\mathcal{P}}\beta_{\sigma}\sigma$ with Pauli expectation values $\beta_{\sigma}(\rho)=\text{tr}(\rho\sigma)$. The fidelity between $\rho$ and |$\ket{\psi}$ is given by
@@
One can bound the number of Pauli measurements $m$ needed on the quantum computer using the SRE [23]:
@@
Ref. [40] proposed an adapted protocol where one estimates the fidelity not in respect to |$\ket{\psi}$, but in respect to a slightly perturbed state |$\ket{\psi^{\prime}}$ which does not feature Pauli expectation values with small, non-zero magnitudes. This incurs an error
@@
where $F_{\text{STAB}}$ is the stabilizer fidelity. $F_{\text{STAB}}$ is lower bounded by $M_{2}$ as shown in Ref. [16]
@@
As $\Delta F\leq 1$, we can get a non-trivial fidelity estimation using the closest stabilizer state when $M_{2}\leq\ln(4/3)$.
Now, let us use the closest stabilizer state to certify the fidelity of random Hamiltonian evolution after time $t$. For evolved state |$\ket{\psi(t)}$, for small $t$ the closest stabilizer state is |$\ket{\psi(0)}$. We have $M_{2}(t)\approx 4t^{2}$, i.e. we can certify the fidelity with non-trivial error up to time $t\leq\frac{1}{2}\sqrt{\ln(4/3)}$.
While $M_{2}$ is small, note that this is not true for SREs with $\alpha<1$. For example, $M_{1/2}(t)\approx\ln(2t^{2})+N\ln(2)\sim N$.
Commonly, Pauli fidelity certification becomes inefficient when states have a lot of nonstabilizerness [23]. However, we argue that this statement applies strictly only for $\alpha>1$. This is because there are two different aspects of nonstabilizerness: While $\alpha<1$ relates to hardness of Clifford simulation, $\alpha>1$ measures the distance to the closest stabilizer states.
For (approximate) Pauli fidelity estimation, the sampling complexity can be related to the closest stabilizer state. As such, one can approximate the fidelity efficiently as long as the $\alpha=2$ SRE $M_{2}$ is sufficiently small. This holds true even when $\alpha<1$ SREs has become extensive.

### v2

@@ 5 Random Hamiltonian evolution
Finally, as we show in Appendix A, the Pauli spectrum and SRE of the GUE evolution matches closely the dynamics of the random basis evolution of Sec. 4. We also observe that the SREs for both models match. Thus, we argue that the scale-invariant behavior that we shown analytically in Sec. 4 for random bases evolution also emerges for the evolution with random Hamiltonians.
6 Complexity and SRE
We now show that SREs can be related to the complexity of different operational tasks, where the type of task depends on $\alpha$. In particular, we relate $\alpha>1$ to approximate fidelity estimation, while $\alpha<1$ to Clifford simulation complexity.
First, we note that for $\alpha>1$, the SRE is similar to $D_{\text{min}}$, which is the distance to the closest stabilizer state [Haug et al.(2024)Haug, Lee, and Kim]. We find that the previously proven lower bound [Haug and Piroli(2023b)] is indeed tight for random evolutions Eq. (24) which we numerically confirm in Appendix C. Note that numeric evidence shows that SREs for $\alpha>1$ also provide an $N$-independent upper bound to $D_{\text{min}}$ [Haug et al.(2024)Haug, Lee, and Kim]. As such, we argue that SREs with $\alpha>1$ probe the closeness to the nearest stabilizer state.
SREs with $\alpha=2$ have been shown an explicit operational meaning: They give a lower bound on fidelity estimation [Flammia and Liu(2011), Leone et al.(2023)Leone, Oliviero, and Hamma]: Given a target state |$\ket{\psi}$, one can estimate the fidelity with actual state $\rho$ by measuring Pauli expectation values [Flammia and Liu(2011)]. The number of samples $m$ to estimate the fidelity is lower bounded as $m\gtrsim\exp(M_{2}(\ket{\psi}))$, while an upper bound is given by $m\lesssim\exp(M_{0}(\ket{\psi}))$ [Leone et al.(2023)Leone, Oliviero, and Hamma].
In Appendix D, we show an explicit fidelity estimation algorithm whose efficiency is directly given by $M_{2}(\ket{\psi})$: Fist, we note that fidelity estimation works especially well when |$\ket{\psi}$ is a stabilizer state. Now, if |$\ket{\psi}$ is not a stabilizer, but close to stabilizer state, then one can use this stabilizer state as a proxy for fidelity estimation: One estimates the fidelity of $\rho$ with the closest stabilizer state $\text{argmax}_{\ket{\psi_{\text{STAB}}}}|\braket{\psi}{\psi_{\text{STAB}}}|^{2}$. We show that when |$\ket{\psi}$ is close to a stabilizer state, this is a good approximation of the fidelity, where the quality depends on $M_{(}\ket{\psi})$. In particular, we show in Appendix D that this scheme has an error $\Delta F\leq 2\sqrt{1-\exp(-M_{2}(\ket{\psi}))}$ and gives a non-trivial approximation of the fidelity as long as $M_{2}(\ket{\psi})\leq\log(4/3)$.
In contrast, SREs with $\alpha<1$ (especially $\alpha=1/2$) show behavior similar to the log-free robustness of magic LR [Howard and Campbell(2017), Rall et al.(2019)Rall, Liang, Cook, and Kretschmer, Liu and Winter(2022)] or max-relative entropy of magic [Bravyi et al.(2019)Bravyi, Browne, Calpin, Campbell, Gosset, and Howard, Liu and Winter(2022)]. They respectively relate to the negativity of the mixture of stabilizer states, or the number of superpositions of stabilizer states needed to represent a given state. LR has been used to estimate fault-tolerant state preparation complexity and relates to the complexity of Clifford based simulation algorithms. These algorithms simulate quantum circuits as Clifford circuits injected with nonstabilizer gates, where the simulation complexity commonly increase exponentially with the number of nonstabilizer gates [Bravyi et al.(2016)Bravyi, Smith, and Smolin]. In fact, $M_{1/2}$ has been used as a proxy for LR to evaluate simulation complexity of Clifford-based simulation algorithms [Rall et al.(2019)Rall, Liang, Cook, and Kretschmer]. Further, we find that the lower bound $\text{LR}\geq\frac{1}{2}M_{1/2}$ Eq. (8) is tight for random evolution (see Appendix C). Additionally, $M_{0}$ is a lower bound to the stabilizer nullity $\nu\geq M_{0}$, which characterizes the complexity of Clifford-based learning algorithms [Gu et al.(2024c)Gu, Oliviero, and Leone, Grewal et al.(2023)Grewal, Iyer, Kretschmer, and Liang, Hangleiter and Gullans(2023), Leone et al.(2024b)Leone, Oliviero, and Hamma].
SRE dynamics and complexity. Now, we study the complexity of different tasks for doped Clifford circuits and random evolution using SREs.
For Clifford gates injected with T-gates, SREs converge for all $\alpha$ to their maximum $M_{\alpha}\sim N$ at T-gate density $q\sim\text{const}$. This is because each T-gate affects only a discrete subset of the Pauli spectrum. We numerically find that $D_{\text{min}}$ and LR appear to show this behavior as well. This implies that fidelity with the closest stabilizer state and classical cost of simulation with Clifford+T correlate. In particular, for $q\sim 1/N$ and thus $M_{\alpha}=\text{const}$, one can efficiently simulate and learn the state [Grewal et al.(2023)Grewal, Iyer, Kretschmer, and Liang, Hangleiter and Gullans(2023), Leone et al.(2024b)Leone, Oliviero, and Hamma], as well as estimate the fidelity [Leone et al.(2023)Leone, Oliviero, and Hamma]. In contrast, for $q\sim\text{const}$ and thus $M_{\alpha}\sim N$ simulation, learning and fidelity estimation is unlikely to be efficient [Bravyi and Gosset(2016), Bravyi et al.(2016)Bravyi, Smith, and Smolin, Leone et al.(2021)Leone, Oliviero, Zhou, and Hamma].
In contrast, SREs for random evolution shows widely different behavior depending on $\alpha$. This is because random evolution affects all Pauli strings even at short evolution times. For $\alpha>1$, $M_{\alpha>1}$ grows as $\propto t^{2}$ and converges to its maximum at $t_{\text{c},\alpha>1}\sim\sqrt{N}$. For $t=\text{const}$, we have $M_{\alpha>1}=\text{const}$ and $D_{\text{min}}=\text{const}$, i.e. the evolution is close to a stabilizer state in terms of fidelity. Thus, at $t=\text{const}$ one can efficiently certify the fidelity of random evolution (see Appendix D).
$\alpha<1$ shows a completely different behavior, growing extremely fast with $t$: For $\alpha=0$, $M_{0}$ is maximal for any $t>0$, which implies that stabilizer nullity $\nu$ is maximal, rendering known near-Clifford learning algorithms inefficient [Grewal et al.(2023)Grewal, Iyer, Kretschmer, and Liang, Hangleiter and Gullans(2023), Leone et al.(2024b)Leone, Oliviero, and Hamma]. For $10<\alpha<$ 1, SREs saturate rapidly at constant evolution time $t_{\text{c},0<\alpha<1}\approx\frac{1}{2}$. Further, extensive $M_{\alpha<1}\sim N$ is reached already for $t\gtrsim 1/\text{poly}(N)$. This hints that simulating random dynamics to arbitrary precision with Clifford simulation algorithms becomes classically intractable already at $t\gtrsim 1/\text{poly}(N)$.
We summarize the growth of SRE for random Hamiltonian evolution with time $t$, and circuits composed of Clifford unitaries and T-gates with T-gate density $q$ in Fig. 6. In particular, we highlight the extremely fast growth in time $t$ for random evolution for $\alpha<1$, which is not observed for $\alpha>1$ or for Clifford+T circuits for any $\alpha$.
Refer to captionRefer to caption
Figure 6: SRE of random Hamiltonian evolution for $\alpha<1$ increases extremely fast with time $t$. In contrast, for $\alpha>1$, as well as Clifford + T circuits for any $\alpha$, the growth in SRE is much slower. We show SRE density $M_{\alpha}/N$ for random Hamiltonian evolution with Eq. (28) and doped Clifford circuits with Eq. (15) for $N=40$.
Finally, we note that the complexity of tasks related to entanglement tasks can be bounded using the stabilizer nullity [Gu et al.(2024c)Gu, Oliviero, and Leone]. For small stabilizer nullity, one can efficiently do entanglement estimation, distillation and dilution. The same tasks become hard when the stabilizer nullity becomes extensive. For Cliffords doped with a constant number of T-gates we have a constant stabilizer nullity, and thus entanglement-related tasks are easy. In contrast, for random evolution the stabilizer nullity becomes already maximal for exponentially small times, where entanglement-related tasks become hard.
7 Discussion
@@
We have studied $\alpha$-SREs for random Clifford circuits doped with T-gates and random time evolution where we demonstrated the connection of Rényi index $\alpha$ to different aspects of complexity of quantum states.
@@
We find that the SRE converges to the maximum at a critical T-gate density $q_{\text{c},\alpha}$ and time $t_{\text{c},\alpha}$ in the thermodynamic limit. We determine the transition exactly for $\alpha=2$, while for general $\alpha$ we determine the convergence using heuristic models of the Pauli spectrum. For $\alpha=2$, we observe universal behavior around the critical point where the derivative of the SRE can be rescaled onto a single curve for all $N$. This hints that the saturation transition is connected to phase transitions, where universal behavior is commonly found, for example for the transition between different phases of quantum many-body system [Osterloh et al.(2002)Osterloh, Amico, Falci, and Fazio] or at complexity transitions in classical and quantum algorithms [Zhang et al.(2022)Zhang, Sone, and Zhuang].
@@
SREs with $\alpha<1$ and $\alpha>1$ probe two different aspects of nonstabilizernes: We find that SREs with $\alpha<1$ behave similar to the log-free robustness of magic LR [Howard and Campbell(2017), Rall et al.(2019)Rall, Liang, Cook, and Kretschmer, Liu and Winter(2022)] or max-relative entropy of magic [Bravyi et al.(2019)Bravyi, Browne, Calpin, Campbell, Gosset, and Howard, Liu and Winter(2022)]. Roughly, these magic monotones relate to the complexity of approximating a state using a superposition of stabilizer states. In contrast, $\alpha>1$ SREs can be related [Haug et al.(2024)Haug, Lee, and Kim] to the min-relative entropy of magic $D_{\text{min}}$ [Bravyi et al.(2019)Bravyi, Browne, Calpin, Campbell, Gosset, and Howard, Liu and Winter(2022)], which measure the fidelity with the closest stabilizer state.
@@
We find that random Hamiltonian evolution has a fundamental separation in nonstabilizerness complexity: Simulating random evolution using Clifford-based algorithms to arbitrary precision is already hard for very small times $t\gtrsim 1/\text{poly}(N)$. In contrast, up to $t=\text{const}$ one can efficiently certify the fidelity of the evolved state within some fixed error $\Delta F\approx 4t$. We show the algorithm in Appendix D by certifying in respect to a stabilizer approximation of the state. In contrast, for Cliffords doped with T-gates, simulation and certification complexity correlate, and become intractable for the same T-gate density $q\gtrsim\log(N)$.
@@
Finally, for both random evolution and Clifford+T model, the critical time and T-gate density is maximal for $\alpha=2$. This indicates that the $2$-SRE holds a special status. Coincidentally, for $\alpha<2$, the SRE is known not to be a monotone [Haug and Piroli(2023b)], while for $\alpha\geq 2$ it is a pure state monotone [Leone and Bittel(2024)].
@@
Finally, we want to highlight the technical contributions of our work which may be of independent interest: We show heuristically that the Pauli spectrum of random Hamiltonian evolution can be approximated by two distinct peaks. With increasing time, the two peaks shift towards each other and eventually merge. This is exactly when the SRE becomes maximal. At last, we introduce a class of random evolution in Eq. (19), which can be seen as evolution in random Clifford bases. This evolution behaves very similar to random Hamiltonian evolution, where we observe numerically the same Pauli spectrum. It can be expressed as random Clifford circuits combined with small-angle single-qubit rotations. This allows us to compute its $2$-SRE analytically for all times $t$. The random Clifford bases evolution could serve as a model of random evolution with an exact circuit representation.
Acknowledgements
@@ Appendix A GUE evolution and random basis evolution
While we believe that both evolutions show similar behavior for polynomial times, we note that for very long times (on the scale of $t\sim 2^{N/2}$ ) both models likely show different behavior in terms of deep thermalization [Cotler et al.(2017)Cotler, Hunter-Jones, Liu, and Yoshida], as the GUE Hamiltonian evolution conserves energy while the other model does not. It has been noted that the ensemble of GUE evolutions forms an exact $k$-design at polynomial times, however stops being a $k$-design at exponential times. This behavior at long times is attributed to energy conservation of the evolution, which at long times leads to a dephasing due to the energy eigenvalues. For non-energy conserved dynamics this behavior at long times is not expected. However, this difference at exponential times is evident in the $k$-design properties, however it may not be evident in the Pauli spectrum and SRE [Haug et al.(2024)Haug, Lee, and Kim].
@@ Appendix D State certification via Pauli measurements and SREs
A common task is state certification to check whether the prepared state $\rho$ is close to the ideal state |$\ket{\psi}$ that one actually wanted to prepare. For this task, Ref. [Flammia and Liu(2011)] proposed a simple algorithm that only requires to measure Pauli strings of the actual state. First, note that one can decompose any state in terms of its Pauli strings, i.e. $\rho=2^{-N}\sum_{\sigma\in\mathcal{P}}\beta_{\sigma}\sigma$ with Pauli expectation values $\beta_{\sigma}(\rho)=\text{tr}(\rho\sigma)$. The fidelity between $\rho$ and |$\ket{\psi}$ is given by
@@
One can bound the number of Pauli measurements $m$ needed on the quantum computer using the SRE [Leone et al.(2023)Leone, Oliviero, and Hamma]:
@@
Ref. [Flammia and Liu(2011)] proposed an adapted protocol where one estimates the fidelity not in respect to |$\ket{\psi}$, but in respect to a slightly perturbed state |$\ket{\psi^{\prime}}$ which does not feature Pauli expectation values with small, non-zero magnitudes. This incurs an error
@@
where $F_{\text{STAB}}$ is the stabilizer fidelity [Bravyi et al.(2019)Bravyi, Browne, Calpin, Campbell, Gosset, and Howard]. $F_{\text{STAB}}$ is lower bounded by $M_{2}$ as shown in Ref. [Haug et al.(2024)Haug, Lee, and Kim]
@@
There is a unique stabilizer state close to |$\ket{\psi}$ as long as $F_{\text{STAB}}(\ket{\psi})>1/2$. Thus, the approximate certification can give a non-trivial result as long as $M_{2}\leq\ln(4/3)$.
Now, let us use the closest stabilizer state to certify the fidelity of random Hamiltonian evolution after time $t$. For evolved state |$\ket{\psi(t)}$, for small $t$ the closest stabilizer state is |$\ket{\psi(0)}$. We have $M_{2}(t)\approx 4t^{2}$ and thus certification error scaling as as $\Delta F\approx 4t$ Thus, one can certify the fidelity of |$\ket{\psi(t)}$ with non-trivial error up to time $t\leq\frac{1}{4}$.
While $M_{2}$ is small for such $t$, note that this is not true for SREs with $\alpha<1$. For example, $M_{1/2}(t)\approx\ln(2t^{2})+N\ln(2)\sim N$ for any $t\gtrsim 1/\text{poly}(N)$.
Commonly, Pauli fidelity certification becomes inefficient when states have a lot of nonstabilizerness [Leone et al.(2023)Leone, Oliviero, and Hamma]. However, we argue that this statement applies strictly only for $\alpha>1$. This is because there are two different aspects of nonstabilizerness: While $\alpha<1$ relates to hardness of Clifford simulation, $\alpha>1$ measures the distance to the closest stabilizer states.
For (approximate) Pauli fidelity estimation, where one checks the fidelity against the closest stabilizer state, the sampling complexity is related to the closest stabilizer state. As such, one can approximate the fidelity efficiently as long as the $\alpha=2$ SRE $M_{2}$ is sufficiently small. This holds true even when $\alpha<1$ SREs has become extensive.